OpenStudy (anonymous):

X+2/17=x/16 what is x+6/19..?

OpenStudy (anonymous):

please solve

OpenStudy (solomonzelman):

the least common denominator for all of those fractions would be 16*17.

OpenStudy (anonymous):

@ Solpoman dear i got 16x+32=17x.... now what should be done... hmmm

OpenStudy (texaschic101):

x + 2/17 = x/16 --- common denominator is 272...multiply everything by 272 272x + 32 = 17x -- subtract 272x from both sides 32 = 17x - 272x 32 = - 255x -32/255 = x now...x + (6/19) -32/255 + 6/19 -- common denominator is 4845 -608/4845 + 1530/4845 = 922/4845 = 314/1615 or is the problem : (x + 6) / 19 ??

OpenStudy (anonymous):

texaschic dear we want to find x+6/19......

OpenStudy (texaschic101):

(x + 6) / 19 or x + (6/19)

OpenStudy (anonymous):

(x+16)/9

OpenStudy (solomonzelman):

 cross multiply and solve for x.